OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6) x=?

OpenStudy (anonymous):

\[7x-14 + 3x+6 = 5x-30\]

OpenStudy (anonymous):

Do you know about the distributive law?

OpenStudy (anonymous):

7x+3x-5x=-30-6+14